OpenStudy (anonymous):

find the range of f f(x)= 2 sin(3x) where x is between (-pi/6) and (pi/6)

OpenStudy (anonymous):

if \(x=-\frac{\pi}{6}\) then \[3x=3\times (-\frac{\pi}{6})=-\frac{\pi}{2}\]

OpenStudy (anonymous):

\[\sin(-\frac{\pi}{2})=-1\] and that is the smallest that since can be then \[2\sin(-\frac{\pi}{2})=2\times (-1)=-2\] and that is your minimum value for the range

OpenStudy (anonymous):

thanks!!!

OpenStudy (anonymous):

oops i meant to say "that is the smallest that SINE can be"

OpenStudy (anonymous):

hint, you will find the largest value similarly if you replace \(x\) by \(\frac{\pi}{6}\)

OpenStudy (anonymous):

yw